# SemHiTok: A Unified Image Tokenizer via Semantic-Guided Hierarchical Codebook for Multimodal Understanding and Generation

### v4

@@ 5 Conclusion
In this work, we introduce SemHiTok, a unified image tokenizer that implements better trade-off between semantic and pixel information, and is fully compatible with and readily deployable within existing next-token MLLMs architectures. SemHiTok innovatively utilizes semanitc-guied hierarchical codebook (SGHC) to realize the reconstruction capability of pixel features without affecting the understanding performance of the original semantic codebook, and achieves SOTA performance on multimodal understanding under LLaVA-v1.5 setting and on ImageNet-50k reconstruction in unified image tokenizers. We further develop a unified MLLM with SemHiTok, which demonstrates competitive performance on both understanding and generation tasks compared to existing unified MLLMs. This highlights the strong potential of SemHiTok and further bridges the gap between discrete and continuous tokenizers, provide the community with a powerful discrete tokenizer.

### v5

@@ 6 Conclusion
Conclusion: We present SemHiTok, a unified image tokenizer featuring a Semantic-Guided Hierarchical Codebook (SGHC) to balance semantic depth with reconstruction fidelity. SemHiTok achieves SOTA results on LLaVA-v1.5 understanding and ImageNet-50k reconstruction benchmarks. When integrated into a unified MLLM, it delivers competitive performance in both understanding and generation, offering a powerful, discrete, and fully compatible alternative to existing tokenizers.
Limitation: We present two limitations: (1). Low generation efficiency: Owing to the use of standard image quantization methods and settings, each 256-resolution image is represented by 256 tokens. This results in relatively low efficiency and high computational cost. (2). Unified large model potential: In tasks involving natural language understanding and multimodal reasoning, advanced post-training techniques such as Chain-of-Thought (CoT) are not explored. This remains a promising direction for future research in the community.
Future Work: In the future, we can explore the potential of unified image tokenizers and test their performance on more difficult tasks, such as image editing and multiple rounds of conversations. In addition, improving the compression ratio of the model and designing a tokenizer that serializes the image in one dimension are also expected.
Acknowledgement
This work is supported by National Key Research and Development Program of China(2024YFE0203100), Scientific Research Innovation Capability Support Project for Young Faculty (No.ZYGXQNJSKYCXNLZCXM-I28), National Natural Science Foundation of China (NSFC) under Grants No.62476293 and No.62372482, and General Embodied AI Center of Sun Yat-sen University.
Ethics Statement
This research does not involve potentially harmful insights, methodologies, or applications, and it raises no concerns regarding conflicts of interest, sponsorship, discrimination, bias, fairness, privacy, security, legal compliance, or research integrity.
Reproducibility statement
Data. The training datasets employed in this study are detailed in Appendix.7.3 and Appendix. 7.4, and all are publicly accessible open-source resources.
Method. To support reproducibility, we elaborate on the methodological details in Section.2, and the implementation will be open-sourced after acceptance.
Performance. All evaluations are carried out on open benchmarks, thereby ensuring the reproducibility of our results.
Code And Model Weight. We will open-source the code and model weight files.